OpenStudy (anonymous):

Best Painters charges an initial fee of $30 plus $24 per hour. Quality Painters does not have an initial fee and charges $28 per hour. At how many hours is the cost the same for the two companies? please explain.

OpenStudy (bahrom7893):

30+24x = y 28x = y 28x = 30+24x 4x = 30 x = 7.5hrs

OpenStudy (anonymous):

Thank you so much !